# The Role of Privacy Guarantees in Voluntary Donation of Private Health Data for Altruistic Goals

### v2

@@ 4.1.2. Response Distribution and Statistical Analysis Results
We find that privacy expectations differ based on the data-collecting entity. In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, use restriction to a specific purpose, expert auditing, and self-auditing. In contrast, a higher percentage of respondents (43%-76%) expected non-profit entities to employ these mechanisms, except for self-auditing (AG2). For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): ”I also assume that the data is looked at in the aggregate and likely no one at the company knows me”. In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: ”I trust that the organization will uphold strict privacy and ethical standards”.
@@
In the for-profit scenario, all privacy statements, except for the access-control statement (PG2), significantly increase the corresponding privacy expectations ($p$-value ¡ $0.05$). In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration (PG3) significantly increases expectations (from 57% to 100%). Comparing the for-profit and non-profit subsets of the descriptive results (1st and 2nd column of Figure 7), we observe that presenting a privacy statement was particularly beneficial in for-profit scenarios. Indeed, the presence of a privacy statement raised respondents’ privacy expectations to nearly the same level as non-profit scenarios for anonymization (PG1, 68% and 67%) and access control (PG2, 58% and 42%). It also reduced the gap in expectations for data expiration (PG3, 67% and 100%)) and purpose restriction (PG4, 74% and 84%)).
@@
In both scenarios, the auditing statements increase the expectation that a given privacy guarantee will be audited ($p$-value ¡ $0.05$). However, we note that incorporating an audit statement does not markedly change privacy expectations. The statistical significance of the auditing statements in Figure 8 indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection. As observed in  Figure 7, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
4.1.3. Qualitative Analysis of Privacy Expectations via Open-Answer Responses
@@
Satisfaction with the privacy guarantees..Satisfaction with the privacy guarantees.Satisfaction with the privacy guarantees.. 16.3% and 20.8% of respondents in the for-profit and non-profit scenarios, respectively, were convinced by the received privacy statement. Some respondents, like P56, vaguely stated that the policy looks promising: ”I see their policy, and they have to follow their own policy”. Others felt explicitly safer and trusted the statement. For example, P494 noted: ”Trust is very important when it comes to medical data. I believe the organization has privacy policies that outline that they collect and will use my data. I also believe the organization will employ security measures to safeguard data”. Similarly, P555 highlighted the presence of the privacy statement: ”(…) purposely states there is software in place to conserve privacy”. and P516 felt their information was safe because of it: ”(…) with the privacy protection in place, they are isolating the data they need while basically ”throwing out” the rest by putting in under that protection. In essence, my information is safe, and they’re only using what they said they’d use”. P208 also felt safer contributing to research knowing the privacy mechanisms were in place: ”I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above”.
@@
Belief in legal and reputational obligations..Belief in legal and reputational obligations.Belief in legal and reputational obligations.. 16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements and reputational concerns. For example, P28 stated regarding for-profits: ”A for-profit organization wouldn’t want to violate HIPAA, HITECH laws”. Additionally, P117 wrote ”I would trust them to do the right thing so they won’t face lawsuits”. Regarding non-profits, P141 noted: ”I trust that the law will restrict any data leaks to third parties”.
@@
Trust in the recipient entity..Trust in the recipient entity.Trust in the recipient entity.. 13.9% and 17.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed a general trust in the recipient entity without specifying reasons. For instance, P70 succinctly said ”I feel they are reliable and trustworthy” and P50 stated that ”I assume they take their research seriously, so they would handle the data carefully”. Other respondents, like P145, reported having no reason not to trust it: ”I have no reason to think they would do anything nefarious with my medical data”.
@@
Trust in the auditing process..Trust in the auditing process.Trust in the auditing process.. 7.8% and 7.9% of respondents in the for-profit and non-profit scenarios, respectively, found the auditing statement to add at least some reliability. For example, P202 showed some reservation but found confidence in the public nature of audits: ”I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public. That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute”. Additionally, P540 had trust in the entity’s data handling because ”they let outsiders audit them” and P51 (who received the expert-auditing statement) because of ”safelocks and checks in place”. On the other hand, one interviewee doubted the expert auditing process and stated: ”somebody else says something doesn’t mean that it’s real”. No survey respondent explicitly raised concern towards the auditing process.
@@
Skepticism on the privacy statement and limits of privacy-preserving guarantees..Skepticism on the privacy statement and limits of privacy-preserving guarantees.Skepticism on the privacy statement and limits of privacy-preserving guarantees.. 21.2% and 10.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed general distrust in the feasibility of the privacy statement. In the non-profit scenario, P34 stated that ”no privacy technology is foolproof”. Similarly, in the for-profit scenario, P45 wrote: ”I don’t trust that the privacy-preserving tech would work.” More precisely, privacy statements were considered too ideal to be fully enforced. Some respondents believed that unauthorized employee access would be unavoidable. For instance, P71 said: ”Why would I trust someone other than my doctor with my medical records? These days especially, I don’t trust anyone. There could be a breach or simply people I don’t know from a hole in the wall will then have access to all my med records. Insane!”. Some others felt that data breaches and cyberattacks were inevitable. For instance, P83 stated that ”the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked”, while P124 noted that ”corporate data breaches are very common”, and P80 echoed this sentiment, saying: ”I believe the [intentions] will be good, but data can be hacked”. We observe that the auditing statement did not substantially instill trust on all accounts — with the exception of expert auditing for access control in non-profit scenario, and data expiration in for-profit scenario. For example, P474 mentioned: ”They can check their privacy technology all they want but when there is a breach it is done and info is stolen. After it fails then they say sorry and offer monitoring but the info is still stolen”.
@@
Doubt on the recipient entity’s motivation to employ PETs..Doubt on the recipient entity’s motivation to employ PETs.Doubt on the recipient entity’s motivation to employ PETs.. 8.6% and 3.3% of the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all. Most criticism was directed at for-profit entities. For instance, P55 noted that for-profit entities ”will do what is profitable and not much more than that”. Similarly, P119 stated ”(…) because it is a for-profit organization. I expect them to cut corners” and P32 wrote that ”for-profit organization have low standard of morality”. However, some respondents also expressed concerns about non-profit entities. For example, P190 in the non-profit scenario remarked: ”Medicine has become a business. My data is only useful to them if it helps them make more money. Money comes first before the actual well being of humans”.
4.2. RQ2: Willingness to donate health data
Next, we answer the second research question (RQ2, see § 1): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?”. As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
4.2.1. Analysis Set Up
In Figure 9, we summarize the respondents’ willingness to donate their health data to a recipient entity. The 1st column (“overall”) reports results for all entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity. The values represent the percentage of respondents willing to donate their heath data to the recipient entity.
We constructed a logistic regression model to understand the factors that influence willingness to donate. We first analyze and confirm the influence of the recipient entity and then separate the participants based on the recipient entities and analyze them in separate models. The results are summarized in Table 3.
@@ 4.2.1. Analysis Set Up
4.2.2. Response Distribution and Statistical Analysis Results
Figure 9. Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario (e.g., the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.) The agreements are binarized here to present the overall tendency.
Figure 10. Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded). The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (310). $N=454$.
@@ 4.2.2. Response Distribution and Statistical Analysis Results
Table 3. Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
@@
Privacy statements and expectations are not the only reason for donation intentions however. While participants were generally less willing to donate to for-profit entities (OR=0.38, $p=\textless 0.001$), when analyzing only control group participants, we observe that the absolute difference in willingness to donate between for-profit and non-profit entities was 8.8% was smaller than the 20.3% difference in privacy expectations between the two entities. This indicates that other non-privacy-related factors, such as the perceived impact or benefits of the donations, also influence participants’ willingness to donate.
Furthermore, we observe that respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.46, $p=0.007$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement. They have heightened privacy concerns that need to be addressed. For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years.”
Finally, we found that prior donors are more willing to donate to both entities. Respondents who have donated before are more willing to donate to both entities, with the OR value of 3.25 ($p=0.011$) and 2.25 ($p=0.033$) for the for and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments. For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes.” and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection”.
4.2.3. Qualitative Analysis of Donation Intention Via Open-Answer Responses
@@
One non-privacy-related reason for donating included supporting research: e.g., P107 expressed that ”there are a number of things I could share to research such as money, time, data, and more. I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines.” Similarly, P97 stated: ”even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist in.”
@@
Ego-centric reasons, particularly familial connections, also motivated donations. P154 mentioned donating because ”my dad died when his heart failed on him, and I’d have rather that not had happened”, and P233 noted that they ”would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did.”
@@
However, their existing (negative) experiences with for-profit healthcare entities also deter some of them from making donations. For example, P136 claimed they have heard a number of times that the entity says they won’t sell your information but can get around it by calling it something else, thus technically selling the information. P244 also mentioned that their information has been leaked before by a clinician.
Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that ”I would give them my data if I got paid for it. I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled.”
@@
An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use. For instance, participants P495, P463, and P459 all share this perception. However, we note that some individual participants alternatively were motivated to donate due to egocentricity, even to for-profit entities. For example, P45 ”just lost my father to heart failure thirteen days ago. (…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.
@@
Respondents expressed negative sentiments toward for-profit entities collecting data for research. The desire for personal reward related closely to lack of donation intent for for-profit organizations. In the for-profit scenario, 11.4% explicitly refused due to negative perceptions of such organizations. For instance, P191 and P202 respectively stated: ”I have a negative connotation with for profit organization” and ”I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me.” Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further. For instance, P36 wrote: ”For profit organization makes profit using the data. I’m never going to donate it. They can BUY it from me for a reasonable amount.”
@@
Alternatively, a small group of 1.3% respondents explicitly mentioned that they were willing to donate data to entities of non-profit nature, but even they expressed concerns about potential misuse or unethical behavior within these organizations. For instance, P122 mentions: ”it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with. Scares me a bit.” Then, P282 notes: ”it is difficult to police everyone in a non-profit organization. I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have.”
@@
Privacy-related donation considerations focused on data sensitivity and leakage. Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that ”there are so many data leaks from so called safe places. No one can anticipate what hackers can do in the future.” Others cited previous data leak experiences as to why they would not donate, as P290 said ”my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her. For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others.” Some respondents who would not donate data also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that ”regardless of privacy policy, the probability of a leak or misuse is high. unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.”.
@@
Another reason was that the participant was concerned that their medical data included too much or too specific information. P155 highlighted this concern: ”I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not”; as did P229: ”It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me.” Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., ”I don’t have mayor illnesses or nothing to hide so Im ok with that”) or because they already share information with other entities (e.g., ”Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different”).
5. Concluding Discussion
Like our survey, when soliciting data donations, recipient entities usually provide an introduction motivating their need to collect the individual’s data, outline privacy guarantees, and seek consent for data donation. Privacy statements function similarly to advertising (O’Flaherty, 2022) by shaping and influencing consumers’ expectations of privacy (Schaub et al., 2015). Yet, these statements do not always have impact.
Prior work has highlighted the ways in which such statements can fail due to being vague or inaccurate. Our work highlights how the impact of privacy statements is moderated by people’s pre-existing privacy expectations.

### v3

@@ IV-A2 Response Distribution and Statistical Analysis Results
We find that privacy expectations differ based on the data-collecting entity. In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, purpose restriction, expert auditing, and self-auditing. In contrast, a higher percentage of respondents (43%-76%) expected non-profit entities to employ these mechanisms, except for self-auditing (AG2). For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): “I also assume that the data is looked at in the aggregate and likely no one at the company knows me”. In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: “I trust that the organization will uphold strict privacy and ethical standards.”
@@
In the for-profit scenario, all privacy statements, except for the access-control statement (PG2), significantly increase the corresponding privacy expectations ($p$-value < $0.05$). In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration (PG3) significantly increases expectations (from 57% to 100%). Comparing the for-profit and non-profit subsets of the descriptive results (1st and 2nd column of Figure 7), we observe that presenting a privacy statement was particularly beneficial in for-profit scenarios. Indeed, the presence of a privacy statement raised respondents’ privacy expectations to nearly the same level as non-profit scenarios for anonymization (PG1, 68% and 67%) and access control (PG2, 58% and 42%). It also reduced the gap in expectations for data expiration (PG3, 67% and 100%)) and purpose restriction (PG4, 74% and 84%)).
@@
Auditing statements significantly increased expectations of corresponding audits in both for-profit and non-profit scenarios ($p$-value < $0.05$). However, we note that incorporating an audit statement does not significantly change privacy expectations. The statistical significance of the auditing statements in Figure 8 indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection. As observed in  Figure 7, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
IV-A3 Qualitative Analysis of Privacy Expectations via Open-Answer Responses
@@
Satisfaction with the privacy guarantees..Satisfaction with the privacy guarantees.Satisfaction with the privacy guarantees.. 16.3% and 20.8% of respondents in the for-profit and non-profit scenarios, respectively, were convinced by the received privacy statement. Some respondents, like P56, vaguely stated that the policy looks promising: “I see their policy, and they have to follow their own policy.” Others felt explicitly safer and trusted the statement. For example, P494 noted: “Trust is very important when it comes to medical data. I believe the organization has privacy policies that outline that they collect and will use my data. I also believe the organization will employ security measures to safeguard data.” Similarly, P555 highlighted the presence of the privacy statement: “(…) purposely states there is software in place to conserve privacy.” and P516 felt their information was safe because of it: “(…) with the privacy protection in place, they are isolating the data they need while basically ‘throwing out’ the rest by putting in under that protection. In essence, my information is safe, and they’re only using what they said they’d use.” P208 also felt safer contributing to research knowing the privacy mechanisms were in place: “I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above.”
@@
Belief in legal or reputational obligations..Belief in legal or reputational obligations.Belief in legal or reputational obligations.. 16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements or reputational concerns. For example, P28 stated regarding for-profits: “A for-profit organization wouldn’t want to violate HIPAA, HITECH laws.” Additionally, P117 wrote “I would trust them to do the right thing so they won’t face lawsuits.” Regarding non-profits, P141 noted: “I trust that the law will restrict any data leaks to third parties.”
@@
Trust in the recipient entity..Trust in the recipient entity.Trust in the recipient entity.. 13.9% and 17.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed a general trust in the recipient entity without specifying reasons. For instance, P70 succinctly said: “I feel they are reliable and trustworthy” and P50 stated that “I assume they take their research seriously, so they would handle the data carefully.” Other respondents, like P145, reported having no reason not to trust it: “I have no reason to think they would do anything nefarious with my medical data.”
@@
Trust in the auditing process..Trust in the auditing process.Trust in the auditing process.. 7.8% and 7.9% of respondents in the for-profit and non-profit scenarios, respectively, found the auditing statement to add at least some reliability. For example, P202 showed some reservation but found confidence in the public nature of audits: “I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public. That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute.” Additionally, P540 had trust in the entity’s data handling because “they let outsiders audit them” and P51 (who received the expert-auditing statement) because of “safelocks and checks in place”. On the other hand, one interviewee doubted the expert auditing process and stated: “somebody else says something doesn’t mean that it’s real”. No survey respondent explicitly raised concern towards the auditing process.
@@
Skepticism on the privacy statement and limits of privacy-preserving guarantees..Skepticism on the privacy statement and limits of privacy-preserving guarantees.Skepticism on the privacy statement and limits of privacy-preserving guarantees.. 21.2% and 10.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed general distrust in the feasibility of the privacy statement. In the non-profit scenario, P34 stated that “no privacy technology is foolproof”. Similarly, in the for-profit scenario, P45 wrote: “I don’t trust that the privacy-preserving tech would work.” More precisely, privacy statements were considered too ideal to be fully enforced. Some respondents believed that unauthorized employee access would be unavoidable. For instance, P71 said: “Why would I trust someone other than my doctor with my medical records? These days especially, I don’t trust anyone. There could be a breach or simply people I don’t know from a hole in the wall will then have access to all my med records. Insane”! Some others felt that data breaches and cyberattacks were inevitable. For instance, P83 stated that “the primary reason for my distrust is due to past news of companies being hacked by people and their data getting leaked”, while P124 noted that “corporate data breaches are very common”, and P80 echoed this sentiment, saying: “I believe the [intentions] will be good, but data can be hacked”. We observe that the auditing statement did not substantially instill trust on all accounts — with the exception of expert auditing for access control in non-profit scenario, and data expiration in for-profit scenario. For example, P474 mentioned: “They can check their privacy technology all they want but when there is a breach it is done and info is stolen. After it fails then they say sorry and offer monitoring but the info is still stolen.”
@@
Doubt on the recipient entity’s motivation to employ PETs..Doubt on the recipient entity’s motivation to employ PETs.Doubt on the recipient entity’s motivation to employ PETs.. 8.6% and 3.3% of participants in the for- and non-profit scenarios, respectively, claimed that these entities are inherently self-serving and lack the motivation to uphold privacy-preserving guarantees or implement such measures at all. Most criticism was directed at for-profit entities. For instance, P55 noted that for-profit entities “will do what is profitable and not much more than that”. Similarly, P119 stated “(…) because it is a for-profit organization. I expect them to cut corners” and P32 wrote that “for-profit organization have low standard of morality”. However, some respondents also expressed concerns about non-profit entities. For example, P190 in the non-profit scenario remarked: “Medicine has become a business. My data is only useful to them if it helps them make more money. Money comes first before the actual well being of humans.”
IV-B RQ2: Willingness to donate health data
Next, we answer the second research question (RQ2, see § I): “How does the deployment of privacy guarantees and auditing influence people’s willingness to donate their personal health data?” As in the previous subsection, we statistically analyze the responses and then report the qualitative rationales collected from the open-text questions.
IV-B1 Analysis Set Up
In Figure 9, we summarize the respondents’ willingness to donate their health data to the presented recipient entity. The 1st column (“overall”) reports results for both entity types, the 2nd column reports the results of the subset with for-profit entity, and the 3rd column reports the results of the subset with non-profit entity. The values represent the percentage of respondents willing to donate their heath data to the recipient entity.
We constructed a logistic regression model to understand the factors that influence willingness to donate. We first analyze and confirm the influence of the recipient entity and then separate the participants based on the recipient entities and analyze them in separate models. The results are summarized in Table III.
@@ IV-B1 Analysis Set Up
IV-B2 Response Distribution and Statistical Analysis Results
Figure 9: Percentage of respondents willing to donate their personal health data to the recipient entity in each scenario (e.g., the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios that with PG(1) presented but no auditing statement were willing to donate.) The agreements are binarized here to present the overall tendency.
Figure 10: Overlap in respondent’s willingness to donate, privacy expectation for the statement they were shown, and trust that the entity would protect their data as described (respondents in the control – no privacy statement – are excluded). The numbers outside each circle summarize the total number of respondents who were e.g., willing to donate their data (310). $N=454$.
TABLE III: Influences of privacy expectations and demographics on willingness to donate of the three participant groups. OR: odd ratio. CI: confidence interval.
@@ IV-B2 Response Distribution and Statistical Analysis Results
Privacy statements and expectations influence donation intentions, but they are not the only factors. While participants were generally less willing to donate to for-profit entities (OR=0.38, $p$-value < $0.001$), other factors are also at play. Looking at control group participants (who received no privacy statements), we found that the difference in willingness to donate between for-profit and non-profit entities was only 8.8%, despite a much larger 20.3% difference in privacy expectations between these entities. This discrepancy suggests that non-privacy-related factors, such as perceived donation benefits, also influence participants’ willingness to donate.
Furthermore, we observe that respondents with bachelor’s degrees are less willing to donate to a non-profit entity (OR = 0.46, $p$-value = $0.007$), which may be related to their greater awareness of the complexities and potential vulnerabilities with privacy statement. They have heightened privacy concerns that need to be addressed. For example, P42 wanted to verify that the PET worked visually; P97 thought that only the stored data was encrypted, but the data transaction was not protected the same way; and P236 argued that: “privacy-preserving technology that works today may not work in a few years”.
Finally, we found that prior donors are more willing to donate to both entities. Respondents who have donated before were more willing to donate to both entities, with the OR value of 3.25 ($p$-value = $0.011$) and 2.25 ($p$-value = $0.033$) for the for- and non-profit entities, respectively.These respondents expressed strong willingness of supporting science and recognized the importance of medical data in developing new treatments. For example, P40 claimed: “I feel that it is important information to share in hopes that they can find better ways to deal with diabetes”. and P208 responded: “Breakthroughs in science and moving forward in knowledge, our greatly benefited by such Data Collection.”
IV-B3 Qualitative Analysis of Donation Intention Via Open-Answer Responses
@@
One non-privacy-related reason for donating included is helping research. For example, P107 wrote that “there are a number of things I could share to research such as money, time, data, and more. I would be willing to share my medical records to contribute to cancer research in hopes that people will be healthier and to enhance research future medicines”. P97 stated: “even if they are for-profit (which I don’t like), any closer we get to developing better treatments and/or a potential cure to cancer is something I’d be willing to assist”.
@@
Ego-centric reasons, particularly familial connections, also motivated donations. P154 mentioned that the motivation was: “my dad died when his heart failed on him, and I’d have rather that not had happened”, and P233 noted that they “would be willing to help to prevent someone else’s mom from dying of a stroke like [theirs] did”.
@@
However, their existing (negative) experiences with for-profit healthcare entities also deter some of them from making donations. For example, P136 claimed they have heard multiple instances where entities promise not to sell information but circumvent this by using different terminology, effectively still selling the data. P244 also mentioned that their information has been leaked by a clinician in the past.
@@
Some respondents desired personal reward for donating and thus were not interested in donating in our scenario: e.g., P92 explained that “I would give them my data if I got paid for it. I would think twice if they want me to donate the data and make no money from it while the researchers will make money off the of research they compiled.” An alternative explanation could be the perceived imbalance in cost and profit between the donor and the recipient entity, where the donor provides data for free while the recipient entity profits from its use. For instance, participants P495, P463, and P459 all share this perception. However, we note that some individual participants alternatively were motivated to donate due to egocentricity, even to for-profit entities. For example, P45 “just lost my father to heart failure thirteen days ago. (…) I would be more than willing to donate my medical records if it helps develop medicines to make hearts function better.”
@@
Respondents expressed negative sentiments toward for-profit entities collecting data. The desire for personal reward related closely to lack of donation intent for for-profit organizations. In the for-profit scenario, 11.4% explicitly refused due to negative perceptions of such organizations. For instance, P191 and P202 respectively stated: “I have a negative connotation with for profit organization” and “I feel like this type of organization already profits significantly off a large number of people (occasionally off of me as well) and as such, I do not want to give them direct permission to profit further off of me.” Furthermore, respondents felt that for-profits already benefit enough and were unwilling to contribute further. For instance, P36 wrote: “For profit organization makes profit using the data. I’m never going to donate it. They can BUY it from me for a reasonable amount.”
@@
Alternatively, a small group of 1.3% respondents explicitly mentioned that they would donate to non-profits but remained concerned about potential data misuse. For instance, P122 mentions: ‘it’s always going to be in the back of my head that there’s a possibility that it’s being sold or taken or used in some way I’m not okay with. Scares me a bit.” Then, P282 notes: ‘it is difficult to police everyone in a non-profit organization. I have seen leaders in organizations act unethically at times, which makes me think that even a well-intended organization cannot fully control the actions of every employee or volunteer they have.”
@@
Privacy-related donation considerations focused on data sensitivity and leakage. Concerns for not donating were mostly about data leakage, with respondents expressing that they 12.9% and 13.9% participants in the for-profit and non-profit scenario, respectively, were protective of their data and hence would not donate at all, as P545 noted that “there are so many data leaks from so called safe places. No one can anticipate what hackers can do in the future.” Others cited previous data leak experiences as to why they would not donate, as P290 said: “my medical data has been breached in the past by a clinician and I had to file a formal grievance against the health care system that employed her. For this reason, because my medical history was abused, I no longer have any trust and will never voluntarily consent to my medical history/information being shared with others.” Some respondents also expressed that they would only possibly donate to specific entities that have already proven trustworthiness to them, as P357 explained that “regardless of privacy policy, the probability of a leak or misuse is high. unless it is an organization I have had personal interaction with, or am very familiar with, it is unlikely I would donate my medical records.”
@@
Some participants expressed concern about the level of detail in medical data. P155 highlighted this concern: “I would not feel comfortable with so much of my personal health information being shared with an organization that is not involved with my direct medical care, regardless of whether there is an advisory board or not”; as did P229: “It’s a lot of detailed information that I’m worried if it somehow gets in the wrong hands, it could reveal a lot of private information about me.” Contrastingly, some respondents felt comfortable sharing data due to a lack of sensitive information (e.g., “I don’t have mayor illnesses or nothing to hide so Im ok with that”) or because they already share information with other entities (e.g., “Considering I already disclose this information to other organizations (for example, data collection on phone, data compiled through search history), other agencies likely have the information on hand already and so another organization having it is no different”).
V Concluding Discussion
Health data donation decisions involve multiple complex factors beyond privacy considerations. Non-privacy factors, such as recipient reputation and perceived societal benefit, often influence willingness to donate. Moreover, the effectiveness of PETs depends on people’s belief on these protections. PETs’ impact on donation willingness diminishes substantially when individuals misunderstand guarantees or doubt implementation integrity.